# Echoes from the Dark: Galaxy Catalog Incompleteness in Standard Siren Cosmology

### v1

@@
Gravitational wave observations can be combined with galaxy catalogs to constrain cosmology and test modified gravity theories using the standard siren method. However, galaxy catalogs are intrinsically incomplete due to observational limitations, potentially leaving host galaxies undetected and thereby weakening constraints or potentially introducing systematic errors. In this work, we present a self-consistent framework to study catalog incompleteness and host weighting effects, implemented in the publicly available CHIMERA pipeline. We obtain joint cosmological and astrophysical population constraints from 100 binary black hole (BBH) events in a LIGO-Virgo-KAGRA O5-like configuration, using spectroscopic galaxy catalogs with varying completeness levels and stellar-mass host weighting schemes. We find percent-level constraints on $H_{0}$ with complete catalogs, reaching precision of $1.6\%$, $1.3\%$, and $0.9\%$ for constant, linear, and quadratic mass weighting, respectively. As completeness decreases, the precision degrades following a sigmoid trend, with a threshold and steepness that increase for stronger weightings. Simultaneously, the correlation between $H_{0}$ and the BBH population mass scale also increases, making results more sensitive to assumptions about the astrophysical population. Remarkably, $2\%$ precision remains achievable when catalogs contain only $50\%$ of the potential host galaxies within the gravitational wave detection horizon, while $1\%$ precision requires host probabilities scaling with stellar mass squared. The results are robust against host weighting mismodeling, even at moderate completeness levels. This work further highlights the importance of spectroscopic galaxy surveys in standard siren cosmology and provides a pathway for developing the science case of future facilities.
@@ 1 Introduction
Gravitational waves (GWs) from compact binary coalescences (CBCs) are maturing as a robust observational probe to constrain cosmology and fundamental physics, offering a new perspective on existing cosmological tensions (Verde et al. 2019; Moresco et al. 2022; Di Valentino et al. 2025). Because GWs carry direct information on the CBC luminosity distance, they can be used as standard distance indicators or standard sirens (Schutz 1986; Holz & Hughes 2005). With the additional source redshift information, it is then possible to constrain cosmology via the luminosity distance-redshift relation. However, the source redshift cannot be obtained directly from current GW data, as it is degenerate with the source masses.
@@
When an electromagnetic counterpart is identified and the host galaxy redshift is measured, a bright siren measurement becomes possible. To date, the only confirmed bright siren is GW170817 (Abbott et al. 2017b), resulting in a Hubble constant measurement at $15\,\%$ precision (Abbott et al. 2017a). The latest release of the Gravitational-Wave Transient Catalog 4 (GWTC-4.0) by the LIGO-Virgo-KAGRA (LVK) Collaboration contains 219 confident CBC detections, the majority of which are binary black holes (BBHs) without electromagnetic counterparts (Abbott et al. 2019, 2021, 2023a, 2025c). It is therefore essential to develop methods to extract the cosmological information without relying on counterpart detection.
@@
These dark siren methods include two main complementary approaches. The first consists in statistically inferring the redshift from catalogs of potential host galaxies within the gravitational wave localization volume (Schutz 1986; Del Pozzo 2012; Chen et al. 2018; Fishbach et al. 2019; Gray et al. 2020; Palmese et al. 2020; Finke et al. 2021; Gair et al. 2023). The second consists in breaking the mass-redshift degeneracy by modeling intrinsic astrophysical properties. This is possible with spectral sirens, sources whose source-frame mass distribution contains features such as peaks, gaps, or changes in slope (Chernoff & Finn 1993; Taylor et al. 2012; Farr et al. 2019; Mastrogiovanni et al. 2021; Mancarella et al. 2022; Ezquiaga & Holz 2022; Mali & Essick 2025). Additional approaches rely on the cross-correlation between the spatial distribution of GWs and galaxies (e.g., Oguri 2016; Mukherjee et al. 2021; Ferri et al. 2025; Pedrotti et al. 2025).
@@
Recent standard siren analyses (Mastrogiovanni et al. 2023; Gray et al. 2023), including the latest GWTC-4.0 cosmology analysis (Abbott et al. 2025a), have advanced beyond treating catalog and spectral siren methods as independent. Instead, they are jointly combined within a hierarchical Bayesian framework, making it possible to robustly marginalize over the uncertainties of the astrophysical and cosmological parameters simultaneously. Several pipelines are publicly available, including gwcosmo (Gray et al. 2020, 2023), icarogw (Mastrogiovanni et al. 2023, 2024), and CHIMERA (Borghi et al. 2024; Tagliazucchi et al. 2025). For further details on the application of GWs to cosmology, see, for example, Palmese & Mastrogiovanni (2025); Pierra & Mastrogiovanni (2025).
@@
While galaxy catalogs provide valuable information, they are intrinsically limited by observational effects that can cause host galaxies to be missed, potentially weakening cosmological constraints (e.g., Gray et al. 2020; Cross-Parkin et al. 2025) and introducing systematic errors (Perna et al. 2024; Hanselman et al. 2025; Alfradique et al. 2025). This incompleteness directly relates to assumptions about the CBC host galaxy population, as GW sources may preferentially reside in certain types of galaxies that are more easily detected and thus better represented in surveys. The link between CBCs and their host galaxies has been increasingly investigated through population synthesis studies (e.g., Artale et al. 2020; Santoliquido et al. 2022) and observational analyses (Vijaykumar et al. 2024), although more BBH events are needed to establish robust observational constraints. As GW data quality improves, accounting for galaxy catalog incompleteness and weighting will become increasingly important. Extending current mock data analyses to model these effects and test for potential systematic errors in future standard siren measurements is a necessary step forward.
@@
In Borghi et al. (2024, hereafter B24), we provided standard siren forecasts for 100 BBHs in LVK O4- and O5-like configurations. We included a complete galaxy catalog with host galaxies having a stellar mass higher than $\log_{10}(M_{\star}/M_{\odot})>10.5$, obtaining percent-level constraints on $H_{0}$ in the case of a galaxy catalog with spectroscopic redshift measurements. A natural extension would be to include all galaxies regardless of stellar mass. However, this approach would implicitly assume that low-mass galaxies with $\log_{10}(M_{\star}/M_{\odot})\approx 6$ have the same probability of hosting BBH mergers as massive $\log_{10}(M_{\star}/M_{\odot})\approx 12$ galaxies. This strong assumption, which extends the BBH hosting probability over several orders of magnitude, requires taking into account observational effects in galaxy surveys.
@@
In this paper, we present a self-consistent framework to study the effects of galaxy catalog incompleteness and host weighting in standard siren analyses and provide updated and more realistic forecasts. This framework can be readily applied to ongoing and upcoming surveys such as Euclid (Euclid Collaboration et al. 2025) and DESI (DESI Collaboration et al. 2024), to inform the science case of future missions such as WST (Mainieri et al. 2024), and for systematic studies in current cosmological analyses (Agarwal et al. 2025). The paper is organized as follows. In Sect. 2, we introduce the statistical framework and its implementation in the CHIMERA code. In Sect. 3, we describe the generation of mock galaxy and GW catalogs. Sect. 4 provides the results of the incompleteness effects in standard siren constraints. Finally, Sect. 5 concludes the paper and discusses future applications of this method.
@@ 2 Method
We consider a population of GW sources, each described by the source-frame parameters $\boldsymbol{\theta}$ (such as redshift $z$, sky position $\hat{\Omega}$, and binary masses $m_{1},\,m_{2}$), which globally follow a population distribution $p_{\rm pop}(\boldsymbol{\theta}\,|\,\boldsymbol{\lambda})$ described by hyperparameters $\boldsymbol{\lambda}$. Given a data set of $N$ independent GW detections $\{\boldsymbol{d}_{i}\}$, the likelihood of $\boldsymbol{\lambda}$ can be obtained with the hierarchical Bayesian formalism (Loredo 2004; Mandel et al. 2019; Vitale et al. 2022):
@@
which corrects for the fraction of detectable events given a target population, with $P_{\rm det}(\boldsymbol{\theta},\boldsymbol{\lambda})\in[0,1]$ being their detection probability. Equation (1) assumes a scale-invariant rate prior, so that the likelihood depends only on the shape of the population distribution, not on the overall rate normalization (Mandel et al. 2019).
@@ 2.1 Population
To model the BBH population, we separate the hyperparameters $\boldsymbol{\lambda}$ into three categories: those that describe the underlying cosmology $\boldsymbol{\lambda}_{\rm c}$, the mass distribution $\boldsymbol{\lambda}_{\rm m}$, and the redshift distribution $\boldsymbol{\lambda}_{\rm z}$. This allows us to write the population model as:
@@
The first term represents the intrinsic mass distribution, typically described using parametric models, although nonparametric approaches are also being studied (e.g, Edelman et al. 2022; Abbott et al. 2023b; Farah et al. 2025). The second term acts as a “redshift prior”, which can be informed by external electromagnetic data, such as galaxy catalogs. Equation (4) assumes that the intrinsic BBH mass distribution does not evolve significantly on cosmological timescales, as supported by current data (see Callister 2024; Abbott et al. 2025b). More generally, we can incorporate dependencies on host galaxy properties $\mathcal{G}$ (such as stellar mass, metallicity, and star formation rate) that may influence BBH formation, evolution, and ultimately their merger rate. In this case, the population prior is marginalized over galaxy properties,
@@ 2.2 Redshift prior
where $p_{\rm gal}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})$ is the probability that a galaxy with properties $\mathcal{G}$ at redshift $z$ and sky location $\hat{\Omega}$ hosts a GW event, and the factor $(1+z)^{-1}$ converts the source-frame time to the detector-frame time.
As a first step, we introduce the case for a complete galaxy catalog that contains all potential host galaxies. In this case, we can define $p_{\rm gal}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})=p_{\rm cat}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})$ and compute this probability as a weighted mean over the contribution of each galaxy $g$ in the catalog:
@@ 2.3 Incompleteness
Figure 1: Effects of galaxy catalog incompleteness within the CHIMERA framework. (a) Sky localization and pixelization (represented with the grid) for a GW event from a stellar mass-weighted MICECATv2 catalog (the underlying blue points). The marker shows the host true position and pixel, from which the distributions in the subsequent panel are computed. (b) Catalog redshift distribution for complete (light blue), $i<22$ cut (blue), and empty catalog (dark blue flat line). The arrow shows the true host redshift. (c) Completion terms accounting for missing galaxies; dark blue shows the background distribution used in standard siren constraints. (d) Resulting galaxy distribution terms, individually normalized for better comparison.
The completeness function is defined in the domain $[0,1]$ and can be computed as:
@@
where $n_{\rm cat}(z)$ is the catalog number density and $n_{\rm theo}(z)$ is the theoretical target number density, both depending on the properties of the host galaxies $\mathcal{G}$. Compared to the galaxy catalog term in Eq. (7), $n_{\rm cat}(z)$ has to be computed over larger volumes to ensure adequate statistics and smooth out local inhomogeneities. We note that this approach, also used in Finke et al. (2021), differs from the one used in gwcosmo and icarogw, where the observed catalog distribution is instead modeled as the integral of a theoretical distribution down a limiting magnitude (Gray et al. 2023; Mastrogiovanni et al. 2024). The key advantage of measuring the catalog density directly is that it enables modeling completeness effects beyond simpler magnitude cuts, such as color-color selection or stellar mass cuts, more flexibly representing the selection functions of current galaxy surveys.
The distribution of galaxies can be described empirically using the parametric formulation introduced by Schechter (1976). It is convenient to express the theoretical number density of potential hosts as:
@@
where $\Phi(M_{\star},z)$ is the Schechter mass function, ${M_{\star}}_{,\rm min}(z)$ and ${M_{\star}}_{,\rm max}(z)$ are the minimum and maximum masses of galaxies hosting BBH mergers, and $w(M_{\star},z)$ accounts for potential host weighting assumptions (see Sect. 3.2). The use of Schechter functions is widespread in recent standard siren analyses, where galaxy luminosity or absolute magnitude is typically used as a weighting factor (e.g., Abbott et al. 2023c; Mastrogiovanni et al. 2023; Turski et al. 2025). Here, we instead weight by stellar mass directly to build a physically consistent framework tied directly to host galaxy properties. Moreover, value-added catalogs for stellar masses are expected to be released for current wide-field surveys such as DESI (DESI Collaboration et al. 2024) and Euclid (Euclid Collaboration et al. 2025). The background probability distribution of potential host galaxies $p_{\rm bkg}$ can then be obtained by multiplying the theoretical galaxy density by the comoving volume element,
@@
In general, this quantity depends on cosmological parameters $\boldsymbol{\lambda}_{\rm c}$, but not on $H_{0}$ as $n_{\rm theo}$ scales as $H_{0}^{3}$ and $\mathrm{d}V_{\rm c}/\mathrm{d}z$ as $H_{0}^{-3}$. Finally, under the assumption that missing galaxies are homogeneously distributed with respect to the background, we can write
@@
This approach, known as homogeneous completion, differs from multiplicative completion, which assumes that missing galaxies trace the distribution of those present in the catalog (Finke et al. 2021), and variance completion, which incorporates galaxy correlation lengths (Dalang & Baker 2024; Leyde et al. 2024, 2025). In this work, we adopt the homogeneous completion method, which is self-consistent with the incompleteness effects simulated in our analysis (see Sect. 3.5).
With all the necessary components in place, the galaxy term corrected for galaxy catalog incompleteness and host weighting becomes:
@@
Figure 2: Stellar mass and redshift distribution of the parent galaxy catalog. Left panel: Stellar mass function in different redshift bins measured from the original MICECATv2 galaxy catalog. The points indicate the peak number density at each redshift bin. Central panel: Theoretical stellar mass function modeled as a double Schechter function evolving in redshift. Right panel: Redshift distribution of the parent galaxy catalog. The black line shows the theoretical distribution obtained from the product of the cosmological volume element dominating at low-$z$ (orange dotted line) and the integral of the Schechter damping the distribution at high-$z$ (purple dotted line). The lower panel shows the percentage difference, highlighting the median (gray line) and the 16-84 percentile range (gray band).

### v2

@@
Gravitational wave observations can be combined with galaxy catalogs to constrain cosmology and test modified gravity theories using the standard siren method. However, galaxy catalogs are intrinsically incomplete due to observational limitations, potentially leaving host galaxies undetected, thereby weakening constraints and potentially introducing systematic errors. In this work, we present a self-consistent framework to study catalog incompleteness and host weighting effects, implemented in the publicly available CHIMERA pipeline. We obtain joint cosmological and astrophysical population constraints from 100 binary black hole (BBH) events in a LIGO-Virgo-KAGRA O5-like configuration using spectroscopic galaxy catalogs with varying completeness levels and stellar-mass host weighting schemes. We find percent-level constraints on $H_{0}$ with complete catalogs, reaching precisions of $1.6\%$, $1.3\%$, and $0.9\%$ for constant, linear, and quadratic mass weighting, respectively. As completeness decreases, the precision degrades following a sigmoid trend, with a threshold and steepness that increase for stronger weightings. Simultaneously, the correlation between $H_{0}$ and the BBH population mass scale increases, making results more sensitive to assumptions about the astrophysical population. Remarkably, $2\%$ precision remains achievable even when catalogs contain only $50\%$ of the potential host galaxies within the gravitational wave detection horizon, while $1\%$ precision requires host probabilities scaling with stellar mass squared. The results are robust against host weighting mismodeling, even at moderate completeness levels. This work further highlights the importance of spectroscopic galaxy surveys in standard siren cosmology and provides a pathway for developing the science case of future facilities.
@@ 1 Introduction
Gravitational waves (GWs) from compact binary coalescences (CBCs) are maturing as a robust observational probe to constrain cosmology and fundamental physics, offering a new perspective on existing cosmological tensions (Verde et al. 2019; Moresco et al. 2022; Valentino et al. 2025). Because GWs carry direct information on the CBC luminosity distance, they can be used as standard distance indicators or “standard sirens” (Schutz 1986; Holz & Hughes 2005). With the additional source redshift information, it is then possible to constrain cosmology via the luminosity distance-redshift relation. However, the source redshift cannot be obtained directly from current GW data, as it is degenerate with the source masses.
@@
When an electromagnetic counterpart is identified and the host galaxy redshift is measured, a “bright siren” measurement becomes possible. To date, the only confirmed case is GW170817 (Abbott et al. 2017b), constraining the Hubble constant to a precision of ${\sim}10\,\%$ (Abbott et al. 2017a; Hotokezaka et al. 2019; Bulla et al. 2022; Palmese et al. 2024). The latest release of the Gravitational-Wave Transient Catalog 4 (GWTC-4.0) by the LIGO-Virgo-KAGRA (LVK) Collaboration contains 219 confident CBC detections, the majority of which are binary black holes (BBHs) without electromagnetic counterparts (Abbott et al. 2019, 2021, 2023a; Abac et al. 2025d). It is therefore essential to develop methods to extract the cosmological information without relying on counterpart detection.
@@
These “dark siren” methods include two main complementary approaches. The first consists of statistically inferring the redshift from catalogs of potential host galaxies within the gravitational wave localization volume (Schutz 1986; Del Pozzo 2012; Chen et al. 2018; Fishbach et al. 2019; Gray et al. 2020; Palmese et al. 2020; Finke et al. 2021; Gair et al. 2023). The second consists of breaking the mass-redshift degeneracy by modeling intrinsic astrophysical properties. This is possible with “spectral sirens,” sources whose source-frame mass distribution contains features such as peaks, gaps, or changes in slope (Chernoff & Finn 1993; Taylor et al. 2012; Farr et al. 2019; Mastrogiovanni et al. 2021; Mancarella et al. 2022; Ezquiaga & Holz 2022; Mali & Essick 2025). Additional approaches rely on the cross-correlation between the spatial distribution of GWs and galaxies (e.g., Oguri 2016; Mukherjee et al. 2021; Ferri et al. 2025; Pedrotti et al. 2025).
@@
Recent standard siren analyses (Mastrogiovanni et al. 2023; Gray et al. 2023), including the latest GWTC-4.0 cosmology analysis (Abac et al. 2025b), have advanced beyond treating catalog and spectral siren methods as independent. Instead, they are jointly combined within a hierarchical Bayesian framework, making it possible to robustly marginalize over the uncertainties of the astrophysical and cosmological parameters simultaneously. Several pipelines are publicly available, including gwcosmo (Gray et al. 2020, 2023), icarogw (Mastrogiovanni et al. 2023, 2024), and CHIMERA (Borghi et al. 2024; Tagliazucchi et al. 2025). For further details on the application of GWs to cosmology, we refer to Palmese & Mastrogiovanni (2025) and Pierra & Mastrogiovanni (2025).
@@
While galaxy catalogs provide valuable information, they are intrinsically limited by observational effects that can cause host galaxies to be missed, potentially weakening cosmological constraints (e.g., Gray et al. 2020; Cross-Parkin et al. 2025) and introducing systematic errors (Perna et al. 2025; Hanselman et al. 2025; Alfradique et al. 2025). This “incompleteness” directly relates to assumptions about the CBC host galaxy population, as GW sources may preferentially reside in certain types of galaxies that are more easily detected and thus better represented in surveys. The link between CBCs and their host galaxies has been increasingly investigated through population synthesis studies (e.g., Artale et al. 2020; Santoliquido et al. 2022) and observational analyses (Vijaykumar et al. 2024), although more BBH events are needed to establish robust observational constraints. As GW data quality improves, accounting for galaxy catalog incompleteness and weighting will become increasingly important. Extending current mock data analyses to model these effects and test for potential systematic errors in future standard siren measurements is a necessary step forward.
@@
In our previous work (Borghi et al. 2024, hereafter B24), we forecasted constraints on $H_{0}$ from 100 BBH mergers detected in an LVK O5-like scenario. We obtained percent-level constraints on $H_{0}$, while assuming a complete galaxy catalog with spectroscopic redshift measurements and considering only galaxies with stellar masses of $\log_{10}(M_{\star}/M_{\odot})>10.5$ as potential hosts. While this approximation was physically motivated by the expectation that BBHs form in more massive galaxies (e.g., Artale et al. 2020; Santoliquido et al. 2022), a more realistic approach requires including the full galaxy population in the analysis. Naively extending to all masses without additional weighting, however, would implicitly assume that low-mass $\log_{10}(M_{\star}/M_{\odot}){\approx}6$ galaxies have the same probability of hosting BBH mergers as massive $\log_{10}(M_{\star}/M_{\odot})\approx 12$ ones. This strong assumption, which extends the BBH hosting probability over several orders of magnitude, requires accounting for BBH host weighting and observational effects in galaxy surveys.
@@
In this paper, we present a self-consistent framework to study the effects of galaxy catalog incompleteness and host weighting in standard siren analyses, we incorporate it into a state-of-the-art analysis pipeline, and we provide updated and more realistic forecasts. This framework can be readily applied to ongoing and upcoming surveys such as Euclid (Euclid Collaboration et al. 2025) and the Dark Energy Spectroscopic Instrument survey (DESI; DESI Collaboration et al. 2024), used to inform the science case of future facilities such as the Wide-Field Spectroscopic Telescope (WST; Mainieri et al. 2024), and implemented in systematic studies for current cosmological analyses (Agarwal et al. 2025). The paper is organized as follows. In Sect. 2, we introduce the statistical framework and its implementation in the CHIMERA code. In Sect. 3, we describe the generation of mock galaxy and GW catalogs. Sect. 4 provides the results of the incompleteness effects in standard siren constraints. Finally, in Sect. 5 we conclude the paper and discuss future applications of this method.
@@ 2 Method
We consider a population of GW sources, each described by the source-frame parameters, $\boldsymbol{\theta}$ (such as redshift $z$, sky position $\hat{\Omega}$, and binary masses $m_{1},\,m_{2}$), which globally follow a population distribution, $p_{\rm pop}(\boldsymbol{\theta}\,|\,\boldsymbol{\lambda})$, described by hyperparameters, $\boldsymbol{\lambda}$. Given a dataset of $N$ independent GW detections, $\{\boldsymbol{d}_{i}\}$, the likelihood of $\boldsymbol{\lambda}$ can be obtained with the hierarchical Bayesian formalism (Loredo 2004; Mandel et al. 2019; Vitale et al. 2022)
@@
which corrects for the fraction of detectable events given a target population, with $P_{\rm det}(\boldsymbol{\theta},\boldsymbol{\lambda})\in[0,1]$ being their detection probability. Equation (1) assumes a scale-invariant rate prior so that the likelihood depends only on the shape of the population distribution, not on the overall rate normalization (Mandel et al. 2019).
@@ 2.1 Population
To model the BBH population, we separate the hyperparameters, $\boldsymbol{\lambda}$, into three categories: those that describe the underlying cosmology, $\boldsymbol{\lambda}_{\rm c}$; the mass distribution, $\boldsymbol{\lambda}_{\rm m}$; and the redshift distribution, $\boldsymbol{\lambda}_{\rm z}$. This allows us to write the population model as
@@
The first term represents the intrinsic mass distribution, typically described using parametric models, although nonparametric approaches are also being studied (e.g, Edelman et al. 2022; Abbott et al. 2023b; Farah et al. 2025). The second term acts as a “redshift prior,” which can be informed by external electromagnetic data, such as galaxy catalogs. Equation (4) assumes that the intrinsic BBH mass distribution does not evolve significantly on cosmological timescales, as supported by current data (see Callister 2024; Abac et al. 2025c). More generally, we can incorporate dependencies on host galaxy properties, $\mathcal{G}$ (such as stellar mass, metallicity, and star formation rate), that may influence BBH formation, evolution, and ultimately their merger rate. In this case, the population prior is marginalized over galaxy properties,
@@ 2.2 Redshift prior
where $p_{\rm gal}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})$ is the probability that a galaxy with properties $\mathcal{G}$ at redshift $z$ and sky location $\hat{\Omega}$ hosts a GW event, $\psi(z\,|\,\boldsymbol{\lambda}_{\rm z})$ is the BBH merger rate density, and the factor $(1+z)^{-1}$ converts the source-frame time to the detector-frame time.
As a first step, we introduce the case for a complete galaxy catalog that contains all potential host galaxies. In this case, we can define $p_{\rm gal}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})=p_{\rm cat}(z,\hat{\Omega},\mathcal{G}\,|\,\boldsymbol{\lambda}_{\rm c})$ and compute this probability as a weighted mean over the contribution of each galaxy, $g$, in the catalog
@@ 2.3 Incompleteness
Figure 1: Effects of galaxy catalog incompleteness within the CHIMERA framework. (a) Sky localization and pixelization (represented with the grid) for a GW event from a stellar mass-weighted MICECATv2 catalog (the underlying blue points). The marker shows the host true position and pixel, from which the distributions in the subsequent panel are computed. (b) Catalog redshift distribution for complete (light blue), $i<22$ cut (blue), and empty catalogs (dark blue flat line). The arrow shows the true host redshift. (c) Completion terms accounting for missing galaxies; dark blue shows the background distribution used in standard siren constraints. (d) Resulting galaxy distribution terms, individually normalized for better comparison.
The completeness function is defined in the domain $[0,1]$ and can be computed as
@@
where $n_{\rm cat}(z)$ is the catalog number density and $n_{\rm theo}(z)$ is the theoretical target number density, both depending on the properties of the host galaxies, $\mathcal{G}$. Compared to the galaxy catalog term in Eq. (7), $n_{\rm cat}(z)$ has to be computed over larger volumes to ensure adequate statistics and smooth out local inhomogeneities. We note that this approach, also used in Finke et al. (2021), differs from the one used in gwcosmo and icarogw, where the observed catalog distribution is instead modeled as the integral of a theoretical distribution down a limiting magnitude (Gray et al. 2023; Mastrogiovanni et al. 2024). The key advantage of measuring the catalog density with a more physically motivated method is that it enables modeling completeness effects beyond simpler magnitude cuts, such as color-color selection or stellar mass cuts, more flexibly representing the selection functions of current galaxy surveys.
The distribution of galaxies can be described empirically using the parametric formulation introduced by Schechter (1976). It is convenient to express the theoretical number density of potential hosts as
@@
where $\Phi(M_{\star},z)$ is the Schechter mass function, ${M_{\star}}_{,\rm min}(z)$ and ${M_{\star}}_{,\rm max}(z)$ are the minimum and maximum masses of galaxies hosting BBH mergers, and $w(M_{\star},z)$ accounts for potential host weighting assumptions (see Sect. 3.1 and Sect. 3.2). The use of Schechter functions is widespread in recent standard siren analyses, where galaxy luminosity or absolute magnitude is typically used as a weighting factor (e.g., Abbott et al. 2023c; Mastrogiovanni et al. 2023; Turski et al. 2025). Here, we instead weight by stellar mass directly to build a physically consistent framework tied directly to host galaxy properties. Moreover, value-added catalogs for stellar masses are expected to be released for current wide-field surveys such as DESI (DESI Collaboration et al. 2024) and Euclid (Euclid Collaboration et al. 2025). The background probability distribution of potential host galaxies, $p_{\rm bkg}$ can then be obtained by multiplying the theoretical galaxy density by the comoving volume element,
@@
In general, this quantity depends on cosmological parameters, $\boldsymbol{\lambda}_{\rm c}$, but not on $H_{0}$, as $n_{\rm theo}$ scales as $H_{0}^{3}$ and $\mathrm{d}V_{\rm c}/\mathrm{d}z$ as $H_{0}^{-3}$. Finally, under the assumption that missing galaxies are homogeneously distributed with respect to the background, we can write
@@
This approach, known as “homogeneous completion,” differs from “multiplicative completion,” which assumes that missing galaxies trace the distribution of those present in the catalog (Finke et al. 2021), and “variance completion,” which incorporates galaxy correlation lengths (Dalang & Baker 2024; Leyde et al. 2024, 2025). In this work, we adopt the homogeneous completion method, which is self-consistent with the incompleteness effects simulated in our analysis (see Sect. 3.5). With all the necessary components in place, the galaxy term corrected for galaxy catalog incompleteness and host weighting becomes
@@
Figure 2: Stellar mass and redshift distribution of the parent galaxy catalog. Left panel: Stellar mass function in different redshift bins measured from the original MICECATv2 galaxy catalog. The points indicate the peak number density at each redshift bin. Central panel: Theoretical stellar mass function modeled as a double Schechter function evolving in redshift. Right panel: Redshift distribution of the parent galaxy catalog. The black line shows the theoretical distribution obtained from the product of the cosmological volume element dominating at low-$z$ (dotted orange line) and the integral of the Schechter damping the distribution at high-$z$ (dotted purple line). The bottom panel shows the percentage difference, highlighting the median (gray line) and the 16-84 percentile range (gray band).